OpenStudy (swaqqout_kid1):

A store ensures that it has exactly 325 movies on its shelves at the beginning of every week. According to their sales records, the store has sold a minimum of 15 movies per day. If 6 days have passed since the beginning of the week, and the sales records accurately predict results for the number of movies sold, what is the highest number of movies the store could currently have on its shelves? 340 90 235 250 @Ria23

OpenStudy (texaschic101):

325 - (15 * 6) = 325 - 90 = 235